. Two fixed charges, +1.0 x 10 C and -3.0 x 106 C, are placed 10 cm apart. (a) Where maya third charge be located so that the net force acting on this charge is zero? (b) Is the equilibrium stable or not?
Answer #1

The charge 110-6 is  placed at origin and charge (2 is at a distance of d 10cm to the right side of charge 1.

A third charge must be placed at a distance x to the left side of smaller charge 110-6 , so that force F31 on the third chargeq_3 due to charge q_1  (directed towards left) is equal to the force F_{32} on charge q_3 due to q_2 (directed towards right)

kqiq3 31一kq2q3 F 32

Net force on charge q_3 is zero.

F_{3}=-F_{31}+F_{32}=-rac{kq_1q_3}{x^2}+rac{kq_2q_3}{(10+x)^2}=0

rac{kq_1q_3}{x^2}=rac{kq_2q_3}{(10+x)^2}

rac{q_1}{x^2}=rac{q_2}{(10+x)^2}

{q_1}{(10+x)^2}=q_2,x^2

1*10^{-6}{(10+x)^2}=3*10^{-6},(x)^2

(10+2)2-3 (2)

-2x^2+20x+100=0

x=-3.66,cm

That is the third charge is placed on the negative X-axis at x=-3.66,cm.

(Or) the third charge is placed at a distance of 3.66,cm to the left side of charge q_1 and at a distance of 13.66 cnm to the left side of charge q_2.

(b)

The equilibrium is not stable.